PCl₅(g) dissociates into PCl₃(g) and Cl₂(g) when heated at 250°C and 1 atm. If the density of the gas mixture at equilibrium is 4.4 g/L, the fraction of PCl₅(g) dissociated is?
1
Expert's answer
2020-11-23T06:56:23-0500

n=0.023 moles present in the equilibrium

4.4*208g/mol=0.021Moles of PCL5 at the start

Hence, 1mole net increases for every mole that dissociates

0.0233-0.0211=0.022 moles that dissociated.

The Fraction is 0.104, at two figures, that fraction of the original PCL5 is 1/10


"PV=nRT" "(1atm)(1litre)=n(0.08206L-atm\/mol-K)(523K)"